Solve each equation by completing the square. If necessary round to the nearest hundredth.
y^2+ 10y+8=0

Answer by htmentor(1343)

y^2 + 10y + 8 = 0
y^2 + 10y = -8
To complete the square, we take 1/2 of the middle term (1st order term) and square it.
10/2 = 5; 5^2 = 25
Then we need to add this value to both sides, so as not to alter the original equation.
y^2 + 10y + 25 = -8 + 25
(y+5)^2 = 17
Thus the solutions are:
y+5 = +-sqrt(17) -> y = +-sqrt(17) - 5